# MOHPER: Multi-objective Hyperparameter Optimization Framework for E-commerce Retrieval System

### v1

@@ 3.1.3. Hyperparamter Optimization Framework
While various hyperparameter optimization (HPO) frameworks can be applied to MOHPER, we utilize Optuna version 4.0 (Akiba et al., 2019), enabling the evaluation of diverse samplers and configuration settings. Optuna’s flexibility and adaptability make it well-suited for the rigorous testing of multiple HPO strategies within MOHPER. To address the need for concurrent and parallel processing across multiple training trials, we integrate RayTune [^4] https://docs.ray.io/en/latest/tune/index.html, built on the Ray framework [^5] https://github.com/ray-project/ray, which is designed to support scalable, distributed computing within Python [^6] https://www.python.org/ applications (Liaw et al., 2018). This integration supports high-throughput, parallelized experimentation, enhancing the efficiency of testing various sampler configurations within MOHPER.
@@
  • : Humans manually decide the weights of each model and choose CTR-based implicit relevance feedback (IRF) is added for ranking.
@@ 3.1.4. Samplers
In this study, we evaluate the performance of various hyperparameter optimization samplers to identify the most effective configuration for the proposed multi-objective optimization framework. The samplers under consideration include:
@@
Random Sampler: This sampler employs a uniform random search strategy to explore the hyperparameter space without any prior knowledge, providing a baseline for comparison.
@@
Grid Sampler: The grid search method systematically explores a predefined grid of hyperparameter values, ensuring exhaustive evaluation across a fixed set of possible configurations. While computationally expensive, it is useful for identifying optimal configurations within a constrained search space.
@@
Gaussian Process (GP) Sampler: The GP sampler (Močkus, 1975) employs Gaussian processes (Rasmussen, 2003) to model the objective function and iteratively refine the search space by selecting hyperparameter configurations with high expected improvement. This method is particularly effective when function evaluations are costly and the search space is complex.
@@
NSGA-III Sampler: The NSGA-III sampler (Deb and Jain, 2013) is an evolutionary algorithm designed for MOHPO. It applies a non-dominated sorting approach and maintains a diverse set of solutions to balance multiple objectives efficiently. This sampler is particularly known to be suited for problems with a large number of objectives and aims to achieve a well-distributed set of Pareto-optimal solutions.
@@
Tree-structured Parzen Estimator (TPE) Sampler: The TPE sampler (Bergstra et al., 2011) uses a probabilistic model to guide the search, building a density estimate of the objective function and using it to propose new configurations. This model-based approach is designed to be more efficient than random or grid search, especially in high-dimensional spaces.
Each of these samplers is tested within the context of the multi-objective optimization problem to evaluate their efficacy in tuning hyperparameters for e-commerce retrieval tasks.
Table 4. Comparison of Sampler and Objective Type Performance for SERP and CPLP data.
Experiment Setups on SERP data on CPLP data
Sampler Objective def. type # Samples CTR avg. metrics CTCVR avg. metrics Overall avg. metrics CTR avg. metrics CTCVR avg. metrics Overall avg. metrics
Random - 500 0.6131 0.5982 0.6057 0.6956 0.6811 0.6884
Grid - # cases 0.6157 0.6078 0.6118 0.6954 0.6801 0.6877
NSGA-III Separately optimized 100 0.6130 0.6024 0.6077 0.6974 0.6774 0.6874
NSGA-III Separately optimized 500 0.6198 0.6166 0.6182 0.6988 0.6830 0.6909
NSGA-III Weighted sum-based 100 0.6117 0.6009 0.6063 0.6938 0.6810 0.6874
NSGA-III Weighted sum-based 500 0.6214 0.6167 0.6191 0.6919 0.6841 0.6880
GP Weighted sum-based 100 0.6140 0.6194 0.6167 0.7013 0.6843 0.6928
GP Weighted sum-based 500 0.6210 0.6197 0.6203 0.6952 0.6890 0.6921
TPE Separately optimized 100 0.6179 0.6145 0.6162 0.7010 0.6776 0.6893
TPE Separately optimized 500 0.6259 0.6136 0.6197 0.6942 0.6856 0.6899
TPE Weighted sum-based 100 0.6154 0.6124 0.6139 0.6961 0.6841 0.6901
TPE Weighted sum-based 500 0.6231 0.6176 0.6204 0.6911 0.6877 0.6894
3.1.5. Common Baselines on Offline Evaluation
For offline evaluation, we establish several baseline models commonly used in information retrieval tasks to provide comparative benchmarks for our Multi-Objective Hyperparameter Optimization (MOHPO) framework. These baselines are as follows:
BM25: A robust probabilistic retrieval model, BM25 (Robertson et al., 2009) ranks documents based on term frequency and inverse document frequency, making it a baseline for sparse keyword-based retrieval.
LLM-based Retrieval: This method leverages a large language model (LLM) to generate contextual embeddings, capturing semantic nuances in queries and item descriptions for dense retrieval. For Korean, we employ KoBERT [^7] https://github.com/SKTBrain/KoBERT, a fine-tuned variant of BERT (Devlin, 2018).
CLIP-based Retrieval: Utilizing a pretrained CLIP model (Radford et al., 2021), this method encodes queries and items into a shared embedding space. We implement two retrieval scenarios: query-to-image matching and query-to-item description matching. For Korean-language support, we adopt Jina-clip-v2 (Koukounas et al., 2024) and choose to apply 768 dims embeddings, since it does not harm the performance due to (Kusupati et al., 2022).
Hybrid Search: This approach integrates BM25’s sparse matching with vector-based semantic retrieval to enhance ranking performance by leveraging their complementary strengths. Similar hybrid approaches are explored in (Wang et al., 2021; Ghawi and Pfeffer, 2019).
Relevance Filtering with Popularity-based Ranking (RFPR): This baseline employs a two-stage ranking strategy. First, items are filtered based on query relevance (e.g., category or attribute matching). Then, the filtered set is ranked according to popularity, measured via historical user interaction metrics such as click-through rates, purchase frequencies, and cart additions. RFPR is widely used in e-commerce retrieval systems to prioritize high-engagement products, thereby improving overall relevance and user satisfaction. Additionally, we evaluate hybrid configurations that combine retrieval models with popularity-based ranking.
Each baseline method is applied to the evaluation dataset scenarios, as described in Section 3.1.1, allowing us to compare their effectiveness relative to our MOHPO framework.
3.1.6. Objectives and Metrics
As outlined in Section 2.2, our primary optimization goals are CTCVR and CTR, with auxiliary objectives aimed at capturing both user engagement and relevance between products and items, as well as actual conversion intent. To assess the effectiveness of the Multi-Objective Hyperparameter Optimization (MOHPO) framework, we compare it with baseline single-objective models and various multi-objective configurations under different sampler settings. Offline evaluations are conducted on holdout test datasets, while online A/B testing measures real-time performance within a production environment. For offline evaluations, while monitoring multiple metrics that correspond to training objectives, we mainly present nDCG@20, nDCG@100, and precision@100 to succinctly report results. These metrics serve as reliable indicators of user experience on our OHouse platform, as they closely align with engagement behaviors and satisfaction.
@@ 3.2. Experimental Results and Discussion
To assess the efficacy of the proposed MOHPER framework, we benchmark its performance against several baseline approaches. Specifically, we examine its application in two distinct e-commerce scenarios: retrieval with keyword queries for search engine result pages (SERP) and retrieval based on category IDs for product list pages (CPLP). The experimental setup involves optimizing three primary metrics (nDCG@20, precision@100), and varying relevance targets across the following five multi-objective scenarios: (a) CTR (2 objectives), (b) CTCVR (2 objectives), (c) CTR+ CTCVR (4 objectives), (d) CTR+ CTCVR+ CTCAR (6 objectives). For multi-objective training, we adopt two approaches: (i) separate multi-objective tuning, where each objective is trained individually within a multi-objective framework, and (ii) weighted sum-based multi-objective tuning, which aggregates metrics into a single weighted objective for optimization. To ensure fairness in objective selection, we limit hyperparameter optimization (HPO) methods to 100 samples by default. For a broader comparison, we also evaluate further sampled configurations. Grid search baselines explore 1,024 configurations for SERP and 512 for CPLP.
@@
3.2.2. RQ1b. Performance Difference According to Samplers and Setups
To evaluate the effectiveness of different samplers, we compared their performance across varying sample sizes to determine their suitability for specific optimization scenarios. The experimental setup involved training the same objectives outlined in 3.2.1, while varying the type of samplers and the number of samples. The comparative results are summarized in Table 4. Among the evaluated samplers, the TPE sampler consistently demonstrated superior performance across all cases and proved to be applicable in diverse scenarios. In contrast, the GP sampler excelled in settings with a small number of samples; however, it is inherently limited to single-objective optimization unless multi-objective problems are reformulated into a single objective using weighted summation. The NSGA-III sampler, while applicable to multi-objective scenarios without such reformulation, exhibited reduced effectiveness when the number of samples was constrained. The results indicate that increasing the number of samples generally enhances the performance of samplers, albeit at the cost of higher computational overhead. To achieve a more efficient trade-off between performance and computational cost, we introduced a cumulative training strategy, as detailed in the experiments of 3.2.4. This approach leverages incremental sampling and training to boost performance while maintaining computational efficiency.
@@
Table 5. Performance comparison for retrieval strategies on SERP data
@@
Table 6. Performance comparison across cumulative training stage with ablation analysis of voting in the meta-stage
@@
This section demonstrates the effectiveness of voting and cumulative learning strategies in optimization. The results, presented in 6, indicate that voting mechanisms can improve performance compared to no-voting naive optimization. This approach, consistent with the benefits observed in ensemble methods (Chan and Stolfo, 1995; Bauer and Kohavi, 1999; Dietterich, 2000), reduces variance and enhances generalization. On the other hand, cumulative learning further improves performance by iteratively building upon prior knowledge, which reduces the need for large sample sizes and minimizes redundant noise introduced by samplers. We hypothesize that this approach effectively mitigates redundant information, particularly in samplers like Tree-structured Parzen Estimator (TPE), which does not explicitly account for variance during its optimization process. As training progresses, the average differentiation between high-performing (good) and low-performing (bad) samples diminishes due to convergence. Cumulative learning counters this effect by selectively retaining high-quality samples and filtering out redundant or low-quality ones, thereby restoring differentiation and enabling more efficient exploration of the hyperparameter space. Specifically, for TPE, this refined focus on promising regions enhances its ability to identify optimal configurations, leading to uplift performance. Moreover, an additional advantage of splitting # samples into a series of mini-cumulative stages is improved debuggability, and the simplicity of implementation between stages further enhances the practicality and effectiveness of this approach.
@@
We present empirical results from applying our proposed methods to various e-commerce experiments on OHouse retrieval tasks. The experiments on SERP (Search Engine Results Page) were conducted over two weeks, while those on CPLP (Category Product List Page) spanned three weeks. To isolate the performance of MOHPER, we maintained the original search space from previous setups and focused solely on adjusting service hyperparameters. These adjustments included modulating model/feature weights and refining model/feature selections for retrieval or ranking. Table 7. summarizes the results, focusing on key performance metrics: CTR, CTCAR, CTCVR, and per-query engagement rates (Click/Q, Cart/Q, and Purchase/Q). While CTR on CPLP showed a slight decline, all other metrics exhibited consistent improvements, demonstrating MOHPER’s effectiveness in optimizing retrieval performance across multiple objectives. Notably, these gains were achieved without increasing computational costs, as our approach modified only hyperparameters without introducing additional features. Although we do not provide additional online empirical results for fairness and clarification, we note that subsequent series of experiments expanding the search space were successfully launched following these conditioned successes.
@@ 4. Conclusions
This paper presents MOHPER, a novel multi-objective hyperparameter optimization (HPO) framework tailored to the demands of e-commerce retrieval, specifically optimizing for critical performance metrics such as CTR, CTCAR, and CTCVR, which are pivotal for enhancing user engagement and conversions in e-commerce systems. As the first multi-objective HPO framework applied to retrieval and ranking tasks in this domain, MOHPER effectively tackles the inherent complexity of concurrently optimizing diverse and competing metrics in e-commerce. Through extensive evaluation, we demonstrated MOHPER’s ability to achieve superior stability, convergence, and robustness in multi-objectives optimization across various HPO sampling strategies, achieving significant improvements in CTR, CTCVR, and related objectives (RQ1). The integration of hybrid optimization mechanisms further enhanced MOHPER’s performance, delivering measurable gains in relevance and user engagement, thus advancing multi-objective optimization for e-commerce (RQ2). Additionally, we developed advanced techniques such as meta-configuration voting and cumulative training, which uplift MOHPER’s adaptability by shortening sampling costs (RQ3) and also provide inspection changes between each cumulative stage. Lastly, large-scale A/B testing conducted on OHouse’s dataset demonstrated MOHPER’s scalability and validated its practical utility, solidifying its role as a foundational optimization tool within the platform (RQ4). In future work, extending MOHPER’s capabilities could involve exploring new search spaces for training diverse machine learning algorithms and refining the design of objective functions to incorporate factors such as the freshness and diversification of retrieved items.

### v2

@@ 3.1.3. Hyperparamter Optimization Framework
While various HPO frameworks can be applied to MOHPER, we utilize Optuna (Akiba et al., 2019), enabling the evaluation of diverse samplers and configuration settings. Optuna’s flexibility and adaptability make it well-suited for the rigorous testing of multiple HPO strategies within MOHPER. To address the need for concurrent and parallel processing across multiple training trials, we integrate RayTune [^3] https://docs.ray.io/en/latest/tune/index.html, built on the Ray framework [^4] https://github.com/ray-project/ray, which is designed to support scalable, distributed computing (Liaw et al., 2018). This integration supports high-throughput, parallelized experimentation, enhancing the efficiency of testing various sampler configurations within MOHPER.
@@
  • : Humans manually decide the weights of each model, and choose CTR-based implicit relevance feedback (IRF) to be added for ranking.
@@
3.1.4. Common Baselines
We compare ours against the following:
BM25: Lexical sparse retrieval baseline  (Robertson et al., 2009).
Dense Retrieval (KoBERT, CLIP): Vector-based semantic retrieval from queries to product description (texts or images). For Korean, we employ KoBERT [^5] https://github.com/SKTBrain/KoBERT, a fine-tuned variant of BERT (Devlin, 2018) and Jina-clip-v2 (Koukounas et al., 2024), motivated from (Radford et al., 2021), with 768-dim.
Hybrid Search: This approach integrates BM25’s sparse matching with dense retrieval to enhance ranking performance by leveraging their complementary strengths as (Wang et al., 2021; Ghawi and Pfeffer, 2019).
Relevance Filtering with Popularity-based Ranking (RFPR): This baseline employs a two-stage ranking strategy. First, items are filtered based on query relevance. Then, the filtered set is ranked according to popularity, measured via historical user interaction metrics such as view/wish/cart/purchase frequencies. RFPR is widely used in e-commerce ranking systems to prioritize high-engagement products, thereby improving overall relevance and user satisfaction. We also evaluate hybrid methods that combine baselines with RFPR.
@@ 3.1.5. Samplers
We evaluated various HPO strategies as follows:
@@
  • Random Sampler: This sampler employs a uniform random search strategy to explore the HP space without any prior knowledge, providing a baseline for comparison.
@@
  • Grid Sampler: The grid search method systematically explores a predefined grid of HP values, ensuring exhaustive evaluation across a fixed set of possible configurations.
@@
  • Gaussian Process (GP) Sampler: The GP sampler (Močkus, 1975) employs Gaussian processes (Rasmussen, 2003) to model the objective function and iteratively refine the search space by selecting HP configurations with high expected improvement.
@@
  • NSGA-III Sampler: The NSGA-III sampler (Deb and Jain, 2013) is an evolutionary algorithm designed for MOHPO. It applies a non-dominated sorting approach and maintains a diverse set of solutions to balance multiple objectives efficiently.
@@
  • Tree-structured Parzen Estimator (TPE) Sampler: The TPE sampler (Bergstra et al., 2011) uses a probabilistic model to guide the search, building a density estimate of the objective function and using it to propose new configurations.
Table 4. Computational complexity per optimization round for each strategy.
Method Complexity Remarks
TPE (Weighted-sum Based) $\mathcal{O}(kdn+kQL+n\log n)$ Sampling: $kdn$; Retrieval + Eval: $kQL$; Quantile: $n\log n$
TPE (Separate Multi-objective) $\mathcal{O}(kdnm+kQ(L+m)+nm\log n)$ Sampling: $kdnm$; Retrieval: $kQL$; Eval: $kQm$; Quantile: $nm\log n$
NSGA-III (Weighted-sum Based) $\mathcal{O}(gNd+gNQL)$ Sampling: $gNd$; Retrieval + Eval: $gNQL$
NSGA-III (Separate Multi-objective) $\mathcal{O}(gNd+gNQ(L+m)+gN^{2}m)$ Sampling: $gNd$; Retrieval: $gNQL$; Eval: $gNQm$; Sorting: $gN^{2}m$
GP (Weighted-sum Based) $\mathcal{O}(kdn^{2}+kQL+n^{3})$ Sampling: $kdn^{2}$; Retrieval + Eval: $kQL$; GP train: $n^{3}$
  • - $n$: # of evaluated trials (observations), $d$: dimension of the HP space, $m$: # of objectives, $k$: # of sampled candidates per iteration
  • - For NSGA-III, $g$: # of generations, $N$: Population size per generation, $Q$: # of queries used in train/eval, $L$: Abbreviations of costs in search engines
@@ 3.2. Experimental Results and Discussion
To assess the efficacy of the proposed MOHPER framework, we benchmark its performance against several baseline approaches. Specifically, we examine its application in two distinct e-commerce scenarios: retrieval with keyword queries for search engine result pages (SERP) and retrieval based on category IDs for product list pages (CPLP). The experimental setup involves optimizing three primary metrics (nDCG@20, precision@100) and varying relevance targets across the following five multi-objective scenarios: (a) CTR (2 objectives), (b) CTCVR (2 objectives), (c) CTR+CTCVR (4 objectives), (d) CTR+CTCVR+CTCAR (6 objectives). For multi-objective training, we adopt two approaches: (i) separate multi-objective tuning, where each objective is trained individually within a multi-objective framework, and (ii) weighted sum-based multi-objective tuning, which aggregates metrics into a single weighted objective for optimization. We limit HPO methods to 100 samples by default to ensure fairness. For a broader comparison, we also evaluate further sampled configurations. Grid search baselines explore 1,024 configurations for SERP and 512 for CPLP.
@@
3.2.2. RQ1b. Performance Difference Based on Samplers and Setups
Figure 4. Performance differences across various sampler configurations on the SERP dataset.
[…]
result_serp_sampler_compare
To evaluate the effectiveness of different samplers, we compared their performance across varying sample sizes to determine their suitability. The experimental setup involved training the same objectives outlined in 3.2.1, while varying the type of samplers and the number of samples. The comparative results are summarized in Figure 4., and the computational costs are described in Table 4.
While sampler complexity varies, retrieval and evaluation dominate the computational cost. Thus, sampler choice has a trivial impact on performance, while # of samples, search space, and objective definitions play a more significant role. Among the evaluated samplers, the TPE sampler consistently demonstrated superior performance across all cases and proved to be applicable in diverse scenarios. In contrast, the GP sampler excelled in settings with a small number of samples; however, it is inherently limited to single-objective optimization unless multi-objective problems are reformulated into a single-objective using weighted summation. The NSGA-III sampler, while applicable to multi-objective scenarios without such reformulation, exhibited reduced effectiveness when the number of samples was constrained. The results indicate that increasing the number of samples generally enhances the performance of samplers, albeit at the cost of higher computational overhead. To achieve a more efficient trade-off between performance and computational cost, we introduced a cumulative training strategy, as detailed in the experiments of 3.2.4. This approach leverages incremental sampling and training to boost performance while maintaining computational efficiency.
@@
Table 5. Performance change as retrieval strategies on SERP
@@
Table 6. Performance change across the cumulative training stage with ablation analysis of voting in the meta-stage
@@
This section demonstrates the effectiveness of voting and cumulative learning strategies in optimization. The results, presented in Table 6, indicate that voting mechanisms can improve performance compared to naive optimization. This approach, consistent with the benefits observed in ensemble methods (Chan and Stolfo, 1995; Bauer and Kohavi, 1999; Dietterich, 2000), reduces variance and enhances generalization. However, cumulative learning further improves performance by iteratively building on prior knowledge, which reduces the need for large sample sizes and minimizes redundant noise introduced by samplers. We hypothesize that this approach effectively mitigates redundant information, particularly in samplers such as TPE, which does not explicitly account for variance during its optimization process. As training progresses, the average differentiation between high- and low-performing samples diminishes due to convergence. Cumulative learning counteracts this effect by selectively retaining high-quality samples and filtering out redundant or low-quality ones, thus restoring differentiation and enabling more efficient exploration of the HP space. Specifically, for TPE, this refined focus on promising regions enhances its ability to identify optimal configurations, leading to uplift performance. Moreover, an additional advantage of splitting # samples into a series of mini-cumulative stages is improved debuggability, and the simplicity of implementation between stages further enhances the practicality and effectiveness of this approach.
@@
We present empirical results from applying our proposed methods to various e-commerce experiments on OHouse retrieval tasks. To isolate the performance of MOHPER, we maintained the original search space from previous setups and focused solely on adjusting service hyperparameters. These adjustments included modulating model/feature weights and refining model/feature selections for retrieval or ranking. To ensure robust estimation, we employed cumulative training over three stages with 500 sampled configurations per stage. Each round of HPO was conducted with a single compute node with three parallel worker processes. The total optimization process with this setting required approximately 18 hours per stage (total 2 days) on the SERP domain dataset (comprising 5,000 training queries) and 7 hours per run on the CPLP domain (total 21 hours). Notably, our search space was carefully designed to avoid introducing latency bottlenecks. As such, the deployment of new configurations did not negatively impact the stability or responsiveness of the production system.
The experiments on SERP were conducted over two weeks, while those on CPLP spanned three weeks. Table 7. summarizes the results, focusing on key performance metrics: CTR, CTCAR, CTCVR, and per-query engagement rates (Click/Q, Cart/Q, and Purchase/Q). While CTR on CPLP showed a slight decline, all other metrics exhibited consistent improvements, demonstrating MOHPER’s effectiveness in optimizing retrieval performance across multiple objectives. These gains were achieved without increasing computational costs and introducing additional features. Although we do not provide additional online empirical results for fairness and clarification, we note that subsequent experiments expanding the search space were successfully launched following these conditioned successes.
@@ 4. Conclusions
This paper presents MOHPER, a novel multi-objective HPO framework tailored to the demands of e-commerce retrieval, specifically optimizing for critical performance metrics such as CTR, CTCAR, and CTCVR, which are pivotal for enhancing user engagement and conversions in e-commerce systems. As the first multi-objective HPO framework applied to retrieval and ranking tasks in this domain, MOHPER effectively tackles the inherent complexity of concurrently optimizing diverse and competing metrics in e-commerce. Through extensive evaluation, we demonstrated MOHPER’s ability to achieve superior stability, convergence, and robustness in multi-objectives optimization across various HPO sampling strategies, achieving significant improvements in CTR, CTCVR, and related objectives (RQ1). The integration of hybrid optimization mechanisms further enhanced MOHPER’s performance, delivering measurable gains in relevance and user engagement, thus advancing multi-objective optimization for e-commerce (RQ2). Additionally, we developed advanced techniques such as meta-configuration voting and cumulative training, which uplift MOHPER’s adaptability by shortening sampling costs (RQ3) and also provide inspection changes between each cumulative stage. Lastly, large-scale A/B testing conducted on OHouse’s dataset demonstrated MOHPER’s scalability and validated its practical utility, solidifying its role as a foundational optimization tool within the platform (RQ4).